OpenStudy (anonymous):

Solve the compound inequality and graph the solution on the number line. 8-3x≥7 or 6x+2>26

OpenStudy (danjs):

First one \[8 - 3x \ge 7\]\[8 - 7 \ge 3x\]\[1 \ge 3x\]\[x \le 1/3\]

OpenStudy (danjs):

Second \[6x + 2 \gt 26\]\[6x \gt 24\]\[x \gt4\]

OpenStudy (danjs):

overall \[x \le 1/3 ~~or ~~x \gt 4\]

OpenStudy (anonymous):

ok what would it look like on a number ine

OpenStudy (danjs):



OpenStudy (anonymous):

ty

OpenStudy (danjs):

welcome